# The Probabilistic Foundations of Surveillance Failure: From False Alerts to Structural Bias

### v1

@@ Abstract
For decades, forensic statisticians have debated whether searching large DNA databases undermines the evidential value of a match. Modern surveillance faces an exponentially harder problem: screening populations across thousands of attributes using threshold rules rather than exact matching. Intuition suggests that requiring many coincidental matches should make false alerts astronomically unlikely. This intuition fails.
Consider a system that monitors 1,000 attributes, each with a 0.5 percent innocent match rate. Matching 15 pre-specified attributes has probability $10^{-35}$, one in 30 decillion, effectively impossible. But operational systems require no such specificity. They might flag anyone who matches any 15 of the 1,000. In a city of one million innocent people, this produces about 226 false alerts. A seemingly impossible event becomes all but guaranteed. This is not an implementation flaw but a mathematical consequence of high-dimensional screening.
We identify fundamental probabilistic limits on screening reliability. Systems undergo sharp transitions from reliable to unreliable with small increases in data scale, a fragility worsened by data growth and correlations. As data accumulate and correlation collapses effective dimensionality, systems enter regimes where alerts lose evidential value even when individual coincidences remain vanishingly rare. This framework reframes the DNA database controversy as a shift between operational regimes. Unequal surveillance exposures magnify failure, making “structural bias” mathematically inevitable. These limits are structural: beyond a critical scale, failure cannot be prevented through threshold adjustment or algorithmic refinement.
@@ 1.1 From Database Search to Threshold Screening
For over two decades, forensic statisticians have debated a fundamental question: does searching a DNA database of a million profiles weaken the evidential value of a match? Stockmarr (1999) argued that database searching increases coincidental match probability, while Balding (2002) maintained that likelihood ratios preserve evidential weight regardless of search method. This controversy (concerning a single database with exact profile matching) remains unresolved (Storvik, 2006; Kaye, 2013).
@@
Consider a concrete example that reveals the nature of this failure. A surveillance system monitors $k=1000$ binary indicators, each with innocent match probability $p=0.005$ (half a percent chance), and flags individuals matching at least $m=15$ attributes. At first glance, this threshold seems extraordinarily conservative. Requiring an exact match across 15 specific predetermined attributes would have probability $(0.005)^{15}\approx 3\times 10^{-35}$, one in 30 decillion). Such exact-match odds suggest false alerts should be astronomically impossible, even across populations of billions.
But operational systems do not require exact matches across 15 predetermined attributes. Instead, they flag individuals matching at least 15 of any 1000 attributes. This seemingly subtle distinction changes everything. With $\lambda=kp=1000\times 0.005=5$ expected chance matches per innocent person and threshold $m=15$ (three times the expected value), direct calculation gives a per-person false alert probability of
@@
This contrast between the DNA setting and modern multi-attribute surveillance also clarifies why the long-standing DNA database controversy has persisted. Classical DNA searches operate in an extremely sparse false-match regime, where random matches are so rare that enlarging the database does not meaningfully dilute the evidential value of a hit. Multi-attribute threshold screening, however, operates in a dense false-match regime where accumulated coincidences quickly overwhelm the signal. These are mathematically distinct operating regimes; recognizing this distinction resolves the apparent conflict in the DNA literature and motivates the broader probabilistic limits developed in the rest of this paper.
@@ 1.2 Main Contributions
    Unified Bayesian-frequentist view: Both regimes exhibit the same exponential scaling $\exp(\lambda D(c\|1))$, but Bayesian actionability requires $n\lesssim\frac{1-\alpha}{\alpha}\,rs\sqrt{\lambda}e^{\lambda D}$ (for desired PPV $\geq\alpha$) while frequentist reliability allows $n\lesssim\sqrt{\lambda}e^{\lambda D}$ (Section 5). Posterior probabilities collapse when $nq\gg rs$, resolving the Stockmarr-Balding controversy as a regime transition.
@@
    Structural bias through group dominance: Differential surveillance exposure creates exponential outcome disparities that cannot be eliminated through threshold adjustments (Theorems 4.2 through 4.3), formalizing “structural bias” arguments from the algorithmic fairness literature.
@@
    Effective dimensionality under correlation: Spatiotemporal dependencies reduce effective attribute counts to $k_{\text{eff}}\approx A/(2\pi\xi^{2})$ (spatial) or $k_{\text{eff}}\approx k/(2\tau)$ (temporal), connecting surveillance analysis to multiple-testing corrections in genomics and neuroimaging (Section 6).
All results use elementary probability theory with proofs provided.
1.3 Motivation and Background
Modern surveillance and screening systems collect vast amounts of data about individuals and attempt to identify rare targets of interest: potential terrorists, criminals, fraudsters, or security threats. These systems operate by comparing observed attributes of individuals (biometric features, transaction patterns, travel histories, communication metadata) against profiles or watchlists, flagging individuals whose patterns match suspiciously well (Lyon, 2003; O’Neil, 2016; Schneier, 2015).
The fundamental challenge is the curse of dimensionality in coincidence detection: as monitored attributes grow, innocent individuals increasingly match multiple criteria by chance. When populations are large (millions to billions), even extremely rare coincidences become statistically certain, generating overwhelming false alerts. This parallels challenges in cybersecurity (Axelsson, 2000), where intrusion detection systems with 99.9% accuracy produce unmanageable false alarms when monitoring billions of network flows, and in open-set recognition (Scheirer et al., 2013), where classifiers must distinguish known classes from unknown inputs.
1.4 Related Work and Connections
Our analysis builds on classical probability theory (Poisson approximation, Chernoff bounds, concentration inequalities) applied to modern surveillance contexts. The system-level bound
$\Pr(\text{false\allowbreak\ alert})\approx 1\allowbreak-\allowbreak(1-q)^{n}$
corresponds to family-wise error rate (FWER) control in multiple testing (Benjamini and Hochberg, 1995; Storey, 2002), though our large-deviation analysis shows when such control becomes infeasible. The base-rate analysis resolves the Stockmarr-Balding debate (Stockmarr, 1999; Balding, 2002) in forensic DNA statistics by identifying distinct asymptotic regimes. Connections to anomaly detection (Axelsson, 2000), open-set recognition (Scheirer et al., 2013), and algorithmic fairness (Hardt et al., 2016; Chouldechova, 2017; Corbett-Davies et al., 2017; Barocas et al., 2019) appear throughout, as false-positive explosion is a fundamental challenge across screening domains.
@@ 1.5 The Basic Screening Model
Now assuming a general model for $p$, let $X_{i}$ denote the number of matching attributes for individual $i$. Under the null hypothesis (individual $i$ is innocent) and the modeling assumptions detailed in Section 1.6, we have:
@@
The screening system flags individual $i$ as suspicious if they match some unsolved crime on $m$ attributes with: $X_{i}\geq m$ for an integer threshold $m$ (we set $m=\lceil c\lambda\rceil$ later; rounding affects only constant factors). The per-person false alert probability is:
@@ 1.6 Modeling Assumptions and Regime of Validity
    Independence across individuals: The match indicators for different individuals are statistically independent (Remark 2.2). In practice, common-mode events (e.g., mass gatherings, natural disasters) may introduce positive dependence, which would only increase false alert rates beyond our bounds.
@@
    Homogeneous match probability (baseline model): For analytical tractability, we initially assume all attributes have the same match probability $p$. Section 4 extends to heterogeneous populations with different exposure rates $p_{g}$ across demographic groups. Remark 2.1 addresses heterogeneity across attributes within individuals.
@@
    Null hypothesis analysis: We analyze false alerts under the assumption that all $n$ individuals are innocent (the null hypothesis). Detection of actual targets requires separate treatment via signal detection theory and receiver operating characteristic (ROC) analysis (Fawcett, 2006), which is not addressed here. Our results characterize the false positive rate; in practice, system designers must balance this against detection power for true targets.
These assumptions define the scope of our analysis. The core results (Sections 2 through 3) establish fundamental limits under these idealized conditions. Later sections extend the framework: Section 4 analyzes heterogeneous populations with differential exposure, and Section 6 incorporates spatiotemporal correlation structures that reduce the effective dimensionality of attribute spaces.
The large-$k$, small-$p$ regime is particularly natural for modern surveillance systems, where technological advances enable collection of thousands of binary indicators (GPS pings, transaction flags, communication events), but each individual match remains rare. For instance, with $k=1000$ location cells and $p=0.005$ (half-percent chance of presence at any flagged location), we have $\lambda=5$ expected false matches, a moderate value that falls squarely within our analytical framework.
2 Fundamental Probabilistic Limits
2.1 The Poisson Approximation
When $k$ is large, $p$ is small, and $\lambda=kp$ is fixed, the $\mathrm{Binomial}(k,p)$ distribution can be approximated by $\mathrm{Poisson}(\lambda)$ (Feller, 1968). Specifically, if $X\sim\mathrm{Binomial}(k,p)$, then
@@ 2.1 The Poisson Approximation
Remark 2.1 (Heterogeneous Attributes).
We assume homogeneous match probability $p$ across all attributes for tractability. By Le Cam’s theorem (Barbour et al., 1992), when attributes have varying probabilities $p_{1},\ldots,p_{k}$ with $\max_{i}p_{i}\to 0$ and $\lambda=\sum_{i=1}^{k}p_{i}$ fixed, the Poisson approximation $d_{\mathrm{TV}}(X,\text{Poisson}(\lambda))\leq 2\sum_{i=1}^{k}p_{i}^{2}$ ensures our results depend only on $\lambda$, not individual $p_{i}$ values.
@@
2.2 Tail Bounds for the Poisson Distribution
To understand how $q$ depends on $m$ and $\lambda$, we use Chernoff bounds and related concentration inequalities (Hoeffding, 1963; Mitzenmacher and Upfal, 2005) for Poisson random variables.
@@ 2.2 Tail Bounds for the Poisson Distribution
where $D(\alpha\|1)=\alpha\log\alpha-\alpha+1$ is the Poisson rate function (also called the Cramér rate function or Cramér transform of the Poisson distribution with unit mean). This is the Cramér rate function for Poisson random variables, which differs from the binary Kullback-Leibler divergence $D_{\text{KL}}(p\|q)=p\log(p/q)+(1-p)\log((1-p)/(1-q))$ used for Bernoulli random variables. The large-deviation rate is asymptotically exact for real $m/\lambda>1$: the tail probability satisfies $q\asymp\lambda^{-1/2}e^{-\lambda D(m/\lambda\|1)}$ with factorial discreteness contributing only the $\lambda^{-1/2}$ subexponential prefactor. We write $D(c\|1)$ or simply $D$ for brevity when the argument is clear from context.
@@
2.3 Critical Population Size
@@
Proof.
Take $m=\lceil c\lambda\rceil$ with $c>1$; replacing $m$ by $c\lambda$ (ignoring rounding) affects only constants in the bounds. The upper bound in (2) is Lemma 2.3 with $m=c\lambda$: $q\leq\exp(-\lambda D(c\|1))$.
The lower bound follows from Robbins’ refinement of Stirling’s formula $m!<\sqrt{2\pi m}\left(\frac{m}{e}\right)^{m}e^{\frac{1}{12m}}$ (Robbins, 1955).
$\Pr(Y=m)=e^{-\lambda}\frac{\lambda^{m}}{m!}\;\geq\;\frac{1}{\sqrt{2\pi m}}\,\exp\!\Big(-\lambda+m\log\lambda-m\log m+m-\tfrac{1}{12m}\Big).$
Setting $m=c\lambda$ gives
$\Pr(Y=m)\;\geq\;\frac{1}{\sqrt{2\pi c\lambda}}\,\exp\!\Big(-\lambda D(c\|1)-\tfrac{1}{12c\lambda}\Big),$
since
$\lambda+m\log(\lambda/m)+m=-\lambda\big[c\log c-c+1\big]=-\lambda D(c\|1).$
Since $q=\Pr(Y\geq m)\geq\Pr(Y=m)$, the same lower bound applies to $q$.
For the system-level probability, independence across $n$ individuals gives $\Pr(\text{no false alert})=(1-q)^{n}$ and thus $\Pr(\text{false alert})=1-(1-q)^{n}$.
Since $e^{t}\geq 1+t$ for all $t\in\mathbb{R}$, substituting $t=\frac{x}{1-x}$ (valid for $x\in[0,1)$) gives
$1-x=\frac{1}{1+t}\geq e^{-t}=e^{-x/(1-x)},$
and thus (by $\ln(1-x)\leq-x$ and $\ln(1-x)\geq-x/(1-x)$ for $x\in(0,1)$),
$e^{-q/(1-q)}\leq 1-q\leq e^{-q},$
so on the system level:
$1-e^{-nq}\leq 1-(1-q)^{n}\leq 1-e^{-nq/(1-q)}.$
Together with the bounds on $q$ above yields the bounds in (3). The factor 1/(1q)1/(1-q) in the upper bound arises from the inequality $(1-q)^{n}\geq e^{-nq/(1-q)}$; since $q\leq e^{-\lambda D}$, we have $q/(1-q)\leq e^{-\lambda D}/(1-e^{-\lambda D})$, which gives the stated upper bound. When $q\leq e^{-\lambda D}$ is small, $1-q\geq 1-e^{-\lambda D}$ provides a uniform lower bound independent of $n$.
The system is unreliable when the exponent in the lower bound becomes $O(1)$ (that is, when the expected number of false alerts per batch is on the order of one):
$\frac{n}{\sqrt{2\pi c\lambda}}\,e^{-\lambda D(c\|1)-\tfrac{1}{12c\lambda}}\;\approx\;1.$
Solving for $n$ gives
$n_{\mathrm{crit}}\;\approx\;\sqrt{2\pi c\lambda}\;e^{\,\lambda D(c\|1)+\tfrac{1}{12c\lambda}}.$
The $\tfrac{1}{12c\lambda}$ term is tiny for large $\lambda$, so asymptotically we write
$n_{\mathrm{crit}}\;\asymp\;\sqrt{\lambda}\,e^{\,\lambda D(c\|1)},$
or for rough scaling (ignoring the $\sqrt{\lambda}$ factor):
$n_{\text{crit}}\asymp e^{\lambda D(c\|1)}=e^{kp\,D(c\|1)}.$
That is, the critical population size beyond which false alerts almost surely occur. ∎
Remark 2.5 (Numerical Validation).
Monte Carlo simulations (5000 runs per data point) validate the Poisson approximation and system-level probability model: simulated false-alert rates match theoretical predictions with mean absolute error of 0.0015 across the critical transition region (Figure 1(a)). The sharp threshold behavior predicted by Proposition 2.6 is clearly evident in the simulation data.
2.4 Sharp Threshold Behavior

### v2

@@ Abstract
Forensic statisticians have long debated whether searching large DNA databases undermines evidential value of matches. Modern surveillance faces an exponentially harder problem: screening populations across thousands of attributes using threshold rules. Intuition suggests requiring many coincidental matches should make false alerts astronomically unlikely. This intuition fails. Consider a system monitoring 1,000 attributes, each with 0.5 percent innocent match rate. Matching 15 pre-specified attributes has probability $10^{-35}$, one in 30 decillion, effectively impossible. But operational systems may flag anyone matching any 15 of the 1,000. In a city of one million innocents, this produces about 226 false alerts. A seemingly impossible event becomes guaranteed. This is a mathematical consequence of high-dimensional screening, not implementation failure. We identify fundamental probabilistic limits on screening reliability. Systems undergo sharp transitions from reliable to unreliable with small data scale increases, a fragility worsened by data growth and correlations. As data accumulate and correlation collapses effective dimensionality, systems enter regimes where alerts lose evidential value even when individual coincidences remain vanishingly rare. This framework reframes the DNA database controversy as a regime shift. Unequal surveillance exposures magnify failure, making “structural bias” mathematically inevitable. Beyond a critical scale, failure cannot be prevented through threshold adjustment or algorithmic refinement.
@@ 1.1 From Database Search to Threshold Screening
For over two decades, forensic statisticians have debated a fundamental question: does searching a DNA database of a million profiles weaken the evidential value of a match? Stockmarr [1] argued that database searching increases coincidental match probability, while Balding [2] maintained that likelihood ratios preserve evidential weight regardless of search method. This controversy (concerning a single database with exact profile matching) remains unresolved [3, 4].
@@
Consider a concrete example that reveals the nature of this failure. A surveillance system monitors $k=1000$ binary indicators, each with innocent match probability $p=0.005$ (half a percent chance), and flags individuals matching at least $m=15$ attributes. At first glance, this threshold seems extraordinarily conservative. Requiring an exact match across 15 specific predetermined attributes would have probability $(0.005)^{15}\approx 3\times 10^{-35}$, one in 30 decillion. Such exact-match odds suggest false alerts should be astronomically impossible, even across populations of billions.
But operational systems do not require exact matches across 15 predetermined attributes. Instead, they may flag individuals matching at least 15 of any 1000 attributes. This seemingly subtle distinction changes everything. With $\lambda=kp=1000\times 0.005=5$ expected chance matches per innocent person and threshold $m=15$ (three times the expected value), direct calculation gives a per-person false alert probability of
@@
This contrast between the DNA setting and modern multi-attribute surveillance also clarifies why the long-standing DNA database controversy has persisted. Classical DNA searches operate in an extremely sparse false-match regime, where random matches are so rare that enlarging the database does not meaningfully dilute the evidential value of a hit.
Multi-attribute threshold screening, however, operates in a dense false-match regime where accumulated coincidences quickly overwhelm the signal. These are mathematically distinct operating regimes. Recognizing this distinction resolves the apparent conflict in the DNA literature and motivates the broader probabilistic limits developed in this paper.
1.2 Motivation and Background
Modern surveillance and screening systems collect vast amounts of data about individuals and attempt to identify rare targets of interest, such as potential terrorists, criminals, or fraudsters. These systems compare observed attributes—biometric features, transaction patterns, travel histories, communication metadata—against profiles or watchlists, flagging individuals whose patterns match suspiciously well [5, 6, 7].
The fundamental challenge is the curse of dimensionality in coincidence detection [8]: as monitored attributes grow, innocent individuals increasingly match multiple criteria by chance. When populations are large (millions to billions), even extremely rare coincidences become statistically certain, generating overwhelming false alerts. This parallels challenges in cybersecurity [9, 10, 11, 12], where even highly accurate intrusion detectors are swamped by false alarms at scale, and in open-set recognition [13], where classifiers must reliably reject vast numbers of unfamiliar inputs.
1.3 Related Work and Connections
Our analysis builds on classical probability theory (Poisson approximation, Chernoff bounds, concentration inequalities) applied to modern surveillance contexts, situating this work within the framework of large-scale inference [14]. The system-level bound $\Pr(\text{false alert})\approx 1-(1-q)^{n}$ corresponds to family-wise error rate (FWER) control in multiple testing [15, 16, 17], though our large-deviation analysis shows when such control becomes infeasible.
The base-rate analysis resolves the Stockmarr–Balding debate [1, 2] in forensic DNA statistics by identifying distinct asymptotic regimes. Connections to anomaly detection [9, 12], open-set recognition [13], and algorithmic fairness [18, 19, 20, 21, 22] appear throughout: false-positive explosion is a fundamental challenge across all screening domains.
@@ 1.4 Main Contributions
    Unified Bayesian-frequentist view: Both regimes exhibit the same exponential scaling $\exp(\lambda D(c\|1))$, but Bayesian actionability requires $n\lesssim\frac{1-\alpha}{\alpha}\,rs\sqrt{\lambda}e^{\lambda D}$ (for desired PPV $\geq\alpha$) while frequentist reliability allows $n\lesssim\sqrt{\lambda}e^{\lambda D}$ (Section 4.1). Posterior probabilities collapse when $nq\gg rs$, resolving the Stockmarr-Balding controversy as a regime transition.
@@
    Structural bias through group dominance: Differential surveillance exposure creates exponential outcome disparities that cannot be eliminated through threshold adjustments (Section 3.3), formalizing “structural bias” arguments from the algorithmic fairness literature.
@@
    Effective dimensionality under correlation: Spatiotemporal dependencies reduce effective attribute counts to $k_{\mathrm{eff}}\approx A/(2\pi\xi^{2})$ (spatial) or $k_{\mathrm{eff}}\approx k/(2\tau)$ (temporal), connecting surveillance analysis to multiple-testing corrections in genomics and neuroimaging (Section 3.4; technical details in Appendix B).
All main results are proved using elementary probability, with full technical details collected in Appendix A. Appendix C presents illustrative examples based on public datasets (UCI Adult Census and Chicago Crime) that show how the phenomena described here arise with real data.
Notation. The symbols above are defined precisely in later sections: $D(c\|1)$ is the Poisson rate function (Lemma 3.3); $c>1$ is the threshold multiplier with $m=\lceil c\lambda\rceil$ (Section 2.2); $\gamma$ and $k_{0}$ parameterize exponential data growth (Section 3.2); $r$, $s$, and $\alpha$ are Bayesian parameters (Section 4.1); and $k_{\mathrm{eff}}$, $\xi$, $\tau$ quantify correlation effects (Section 3.4). A complete notation summary appears in Section 6.
2 Materials and Methods
@@ 2.1 The Basic Screening Model
More generally, the match probability $p$ may be derived from any domain-appropriate model; the hypergeometric calculation above serves merely as one concrete example. Our analysis holds for any fixed value of $p\ll 1$, regardless of how it is obtained. With this understanding, let $X_{i}$ denote the number of matching attributes for individual $i$. Under the null hypothesis (individual $i$ is innocent) and the modeling assumptions detailed below, we have:
@@
The screening system flags individual $i$ as suspicious if they match some unsolved crime on $m$ attributes with: $X_{i}\geq m$ for an integer threshold $m$. Throughout, we consider thresholds of the form $m=\lceil c\lambda\rceil$ for constant $c>1$, representing detection criteria set above the expected number of chance matches; rounding affects only constant factors. The per-person false alert probability is:
@@ 2.2 Modeling Assumptions and Regime of Validity
    Independence across individuals: The match indicators for different individuals are statistically independent (Remark 3.2). In practice, common-mode events (e.g., large gatherings or seasonal shifts) may introduce positive dependence, which would only increase false alert rates beyond our bounds.
@@
    Homogeneous match probability (baseline model): For analytical tractability, we initially assume all attributes have the same match probability $p$. Section 3.3 extends to heterogeneous populations with different exposure rates $p_{g}$ across demographic groups. Remark 3.1 addresses heterogeneity across attributes within individuals.
@@
    Null hypothesis analysis: We analyze false alerts under the assumption that all $n$ individuals are innocent (the null hypothesis). Detection of actual targets requires separate treatment via signal detection theory and receiver operating characteristic (ROC) analysis [23], which is not addressed here. Our results characterize the false positive rate; in practice, system designers must balance this against detection power for true targets.
These assumptions define the scope of our analysis. The core results (Sections 3.1 through 3.2) establish fundamental limits under these idealized conditions. Later sections extend the framework: Section 3.3 analyzes heterogeneous populations with differential exposure, and Section 3.4 summarizes how spatiotemporal correlation reduces effective dimensionality (with technical details in Appendix B).
The large-$k$, small-$p$ regime is particularly natural for modern surveillance systems, where technological advances enable collection of thousands of attributes per person (GPS pings, transaction records, communication events). In our model, each attribute contributes a binary match indicator (match or no match) of whether it aligns with a profile of interest, and each such match remains rare. For instance, with $k=1000$ location cells and $p=0.005$ (half-percent chance of presence at any flagged location), we have $\lambda=5$ expected false matches, a moderate value that falls squarely within our analytical framework.
3 Results
3.1 Fundamental Probabilistic Limits
3.1.1 The Poisson Approximation
When $k$ is large, $p$ is small, and $\lambda=kp$ is fixed, the $\mathrm{Binomial}(k,p)$ distribution can be approximated by $\mathrm{Poisson}(\lambda)$ [24]. Specifically, if $X\sim\mathrm{Binomial}(k,p)$, then
@@ 3.1.1 The Poisson Approximation
Remark 3.1 (Heterogeneous Attribute Probabilities).
For analytical tractability we assume a homogeneous match probability $p$ across attributes, but this assumption can be relaxed. Let the attribute-level match probabilities be $p_{1},\ldots,p_{k}$, and define
$X=\sum_{i=1}^{k}\mathrm{Bernoulli}(p_{i}),\qquad\lambda=\sum_{i=1}^{k}p_{i}.$
By Le Cam’s theorem [25], the total variation distance between the distribution of $X$ and a Poisson distribution with mean $\lambda$ satisfies
$d_{\mathrm{TV}}\!\left(X,\mathrm{Poisson}(\lambda)\right)\leq 2\sum_{i=1}^{k}p_{i}^{2}.$
Thus the Poisson approximation remains accurate whenever all $p_{i}$ lie in the small-probability regime, even if the $p_{i}$ differ substantially from one another. In this heterogeneous setting, the distribution of $X$, and therefore all results in this section, depends on the collection $\{p_{i}\}$ only through the aggregate rate $\lambda$, up to an approximation error of order $\sum_{i}p_{i}^{2}$.
Group-level heterogeneity (e.g., differing $p_{g}$ across demographic groups) is treated separately in Section 3.3.
@@
Since we screen $n$ individuals independently, the probability that at least one innocent person is flagged is:
$\Pr(\text{at least one false alert})=1-(1-q)^{n}.$
@@
Our analysis assumes independence. Extensions to positively associated arrays via Chen–Stein or Stein’s method [26] are possible but omitted here. The key implication is that positive dependence increases tail probabilities, so our independence-based bounds underestimate true false alert rates when correlation is present. Thus, the independence-based bounds are conservative.
3.1.2 Tail Bounds for the Poisson Distribution
To understand how $q$ depends on $m$ and $\lambda$, we use Chernoff bounds and related concentration inequalities [27, 28] for Poisson random variables.
@@ 3.1.2 Tail Bounds for the Poisson Distribution
where $D(\alpha\|1)=\alpha\log\alpha-\alpha+1$ is the Poisson rate function.
Remark 3.4 (Rate Function Properties).
The function $D(\alpha\|1)$ is also called the Cramér rate function or Cramér transform of the Poisson distribution with unit mean. This differs from the binary Kullback–Leibler divergence $D_{\mathrm{KL}}(p\|q)=p\log(p/q)+(1-p)\log((1-p)/(1-q))$ used for Bernoulli random variables.
The large-deviation rate is asymptotically exact for $m/\lambda>1$: the tail probability satisfies $q\asymp\lambda^{-1/2}e^{-\lambda D(m/\lambda\|1)}$, with factorial discreteness contributing only the $\lambda^{-1/2}$ subexponential prefactor. We write $D(c\|1)$ or simply $D$ for brevity when the argument is clear.
@@
3.1.3 Critical Population Size
@@
Proof sketch.
The upper bound follows directly from Lemma 3.3 with $m=c\lambda$, giving $q\leq\exp(-\lambda D(c\|1))$. The lower bound uses Robbins’ refinement of Stirling’s formula, which shows that $\Pr(Y=m)$ has the same exponential rate $D(c\|1)$ up to a subexponential $\sqrt{\lambda}$ factor. Because $q=\Pr(Y\geq m)\geq\Pr(Y=m)$, these bounds yield matching exponential rates for $q$.
The system-level inequalities (3) follow from $e^{-nx/(1-x)}\leq(1-x)^{n}\leq e^{-nx}$ for $x\in(0,1)$. The critical population scale arises when $nq\sim 1$, giving $n_{\mathrm{crit}}\approx\sqrt{\lambda}\,e^{\lambda D(c\|1)}$, and hence, up to subexponential factors, $n_{\mathrm{crit}}\asymp e^{\lambda D(c\|1)}$ as stated. Full derivations appear in Appendix A. ∎
3.1.4 Sharp Threshold Behavior